A company estimates that 0.3% of their products will fail after the original warranty period but within 2 years of the purchase, with a replacement cost of $150. If they offer a 2 year extended warranty for $8, what is the company's expected value of each warranty sold?
Answer #1

P(company loosing $150 - $8) = 0.003

P(company gaining $8) = 1 - 0.003 = 0.997

Expected value = 8 * 0.997 - (150-8) * 0.003 = $7.55 (ans)